OpenStudy (anonymous):

write the sum using summation notation, assuming the suggested pattern continues. -8 - 3 + 2 + 7 +...+67 @ganeshie8 @dan815 @michele_Laino @nincompoop @solomonzelman @ paki @nnesha @undeadknight26

OpenStudy (michele_laino):

it is an arithmetic sequence, the first term is -8, whereas the last term is 67, furthermore, the constant of the sequence is: -3-(-8)=2-(-3)=...? please complete

OpenStudy (anonymous):

5 @Michele_Laino

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, we have to know how many terms are into your sequence. In order to do that, we have to use this general formula: \[\Large {a_n} = {a_1} + \left( {n - 1} \right)d\]

OpenStudy (michele_laino):

where n is the number of terms, d=5, a_1=-8, and a_n=67

OpenStudy (michele_laino):

substituting those quantities, we can write: \[67 = - 8 + \left( {n - 1} \right) \times 5\]

OpenStudy (michele_laino):

what is n?

OpenStudy (anonymous):

16?

OpenStudy (michele_laino):

that's right! we have n= 16 terms in our sequence

OpenStudy (michele_laino):

so the requested sum S is given by the subsequent formula: \[S = \frac{{{a_1} + {a_n}}}{2} \times n\]

OpenStudy (michele_laino):

where a_1=-8, a_n=67, n=16, please substitute into that formula above

OpenStudy (michele_laino):

\[S = \frac{{{a_1} + {a_n}}}{2} \times n = \frac{{ - 8 + 67}}{2} \times 16 = ...?\]

OpenStudy (anonymous):

472?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

what do i do next?

OpenStudy (michele_laino):

we have completed your exercise

OpenStudy (anonymous):

no i have to put it in sum notation @Michele_Laino

OpenStudy (michele_laino):

then I think that we have to write this: \[\begin{gathered} - 8 + \left( { - 3} \right) + 2 + 7 + 12 + 17 + + 22 + 27 + 32 + 37 + \hfill \\ + 42 + 47 + 52 + 57 + 62 + 67 = 472 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

is it right?

OpenStudy (anonymous):

i need to write it in sum notation form?

OpenStudy (michele_laino):

sorry what do you mean with "notation form"?

OpenStudy (anonymous):

summation notation

OpenStudy (michele_laino):

we can write this: \[\Large \sum\limits_1^{16} { - 8 + \left( {n - 1} \right) \times 5} = 472\]

OpenStudy (michele_laino):

or: \[\Large \sum\limits_1^{16} {\left[ { - 8 + \left( {n - 1} \right) \times 5} \right]} = 472\]

OpenStudy (anonymous):

these are none of my options

OpenStudy (michele_laino):

which can be simplified as follows: \[\Large \sum\limits_1^{16} {\left( {5n - 13} \right)} = 472\]

OpenStudy (michele_laino):

now?

OpenStudy (anonymous):

nope

OpenStudy (michele_laino):

\[\Large \sum\limits_1^{16} n = 136\] now?